OpenStudy (anonymous):

Two lines are perpendicular. The first line passes through (1, –3) and has a y-intercept of (0, 3). Which of the following cannot be the equation of the second line?

OpenStudy (anonymous):

A. y=1/6x-3 B. 6x – y = –2 C. x – 6y = –2 D. 2x = 12y + 3

OpenStudy (anonymous):

Ok so the first step is to find the slope of the first line. They gave you two points so you can find that using the slope formula; \[\frac{ y_{2}-y_{1} }{ x_{2}-x_{1} }\] so: \[\frac{ 3 - (-3) }{ 0-1 }\] that equals -6 so the slope is -6. Now the other line is perpendicular so its slope is the negative reciprocal of the first line so its slope is (1/6). Now solve all of your answer choices for y to get them in form y=mx+b. If one of the lines "m' is not (1/6) thats the answer because then the lines aren't perpendicular because the slopes aren't negative reciprocals of each other.

OpenStudy (anonymous):

ah so it would be d?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no it would not be d i said one which doesn"t have a slope of 1/6 @equqewu4hf

OpenStudy (anonymous):

so its b)